OpenStudy (anonymous):

How do you graph the quadratic function y=(x-1)^2

OpenStudy (anonymous):

plug in numbers for x create a table of values plot points and connect together

OpenStudy (anonymous):

The graph of y=(x-1)^2 is just the graph of y=x^2 shifted one unit to the right.

OpenStudy (anonymous):

Show a picture on the graph